OpenStudy (anonymous):

I have a payroll accounting question: Wilson, a medical sec., earns $2650 monthly for a 37.5 hour week. for overtime work, she recieves extra pay at the regular hourly rate up to 40 hours and time and one-half beyond 40 hours in any week. during one semimonthly pay period, wilson worked 6 hours over-time. only 2 hours of this overtime were beyond 40hours in any week. Compute: a.) the regular semimonthly earning b.) the overtime earning c.) the total earning d.) the overtime rate e.) the regular hourly rate

OpenStudy (anonymous):

a.) the regular semimonthly earning semi means half so how much does Wilson regularly make in half a month?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wilson, a medical sec., earns $2650 monthly (4 weeks) semi-monthly: 2 weeks how much is regular pay?

OpenStudy (anonymous):

we have to compute that

OpenStudy (anonymous):

basically you got: \[\frac{ monthly }{ 4 weeks } = \frac{ semi-monthly }{ 2 weeks }\]

OpenStudy (anonymous):

\[\frac{ $2650 }{ 4 } = \frac{ semi-monthly }{ 2 }\] Answer to a): find x, where "x = regular semi-monthly pay" \[\frac{ $2650 }{ 4 } = \frac{ semi-monthly }{ 2 }\]

OpenStudy (anonymous):

solve for x: regular semi-monthly pay \[\frac{ $2650 }{ 4 } = \frac{ x }{ 2 }\]

OpenStudy (anonymous):

1325

OpenStudy (anonymous):

YES answer to a) $1325

OpenStudy (anonymous):

b.) the overtime earning the overtime earning has two different rates 1. the rate upto 40 hours/wk 2. the rate when exceeding 40 hours/wk

OpenStudy (anonymous):

information: "during one semimonthly pay period, wilson worked 6 hours over-time. only 2 hours of this overtime were beyond 40hours in any week." so question is about the distribution between tariffs 1. and 2.

OpenStudy (anonymous):

regular week: 37.5 hours within extended working time: 4 hours exceeding: 2 hours

OpenStudy (anonymous):

we just need to know the tariffs for 1. and 2., then the earning can be easily calculated with 4 hours, 2 hours

OpenStudy (anonymous):

this also answers questions d) and e).